### v1

Abstract
In this work, we investigate the discrete Calderón problem on grid graphs of dimension three or higher, formed by hypercubic structures. The discrete Calderón problem is concerned with determining whether the discrete Dirichlet-to-Neumann (DtN) operator, which links boundary potentials to boundary current responses, can uniquely identify the conductivity values on the graph edges. We provide an affirmative answer to the question, thereby extending the classical uniqueness result of Curtis and Morrow for two-dimensional square lattices. The proof employs a novel slicing technique that decomposes the problem into lower-dimensional components. Additionally, we support the theoretical finding with numerical experiments that illustrate the effectiveness of the approach.
Keywords: discrete Calderón problem, multi-dimensional grid graphs, discrete Dirichlet-to-Neumann map, conductivity recovery, slicing technique
1 Introduction
The Calderón problem named after Alberto Calderón falls into the class of inverse boundary value problems, mostly investigated in an open bounded domain $\Omega\subseteq\mathbb{R}^{d}$, and focuses on the recovery of the interior conductivity only by the measurements on the boundary $\partial\Omega$. Physically, the interior electric potential follows Kirchhoff’s current law, which takes the form of a second-order elliptic partial differential equation (PDE) in the divergence form. Since the seminal work of Alberto Calderón [12], the problem has been extensively studied in the last fourty years, with many deep theoretical results [5, 27].
The “discrete” Calderón problem focuses on a finite graph $G=(E,D,\partial D)$. The set $E$ contains edges that connect two nodes in the set $D\cup\partial D$. The electric system on the graph can be viewed as connecting homogeneous electric wires according to the graph structure $G$. Then the electric potential function, on both the edges and nodes, that is, the wires and their connecting points, is fully determined by its value on nodes $D\cup\partial D$, and by the linear interpolation of nodal values for each edge.
These two versions of the Calderón problem (continuous versus discrete) represent two fundamentally different ways to formalize the physical problem. In a series of works, Borcea and her collaborators [6, 7, 8, 9, 10] established deep connections between the continuous Calderó problem with the discrete one. However, the mathematical tools utilized to investigate these two versions differ significantly. In particular, very different excitations (i.e., the boundary potential chosen to induce the current measurement) are employed in the analysis. For the continuous Calderón problem, the so-called complex geometric optics solutions [26] with high-frequency oscillations represent one powerful tool to derive theoretical results. In contrast, in the discrete case, the high-frequency property is inherently incompatible with the finite graph structure, and a new type of excitation has been employed, i.e., excitations that are able to localize all interior potential within a sub-domain of the graph [19].
In a graph $G$, $D$ and $\partial D$ are finite sets with no common elements, and $E:=\{pq:p,q\in D\cup\partial D\}$ where $pq$ is an unordered pair. For any node $p\in D\cup\partial D$, its neighbouring nodes $\mathcal{N}(p)$ is defined by the set $\{q:pq\in E\}$. Each edge $pq$ in the graph can also be assigned a positive conductivity value $\gamma$, where by definition it holds that $\gamma_{pq}=\gamma_{qp}>0$. Moreover, real-valued functions on a finite set $U$ can be identified with vectors in $\mathbb{R}^{U}$, in which way we define the potential function (or vector) as $\mathbf{u}\in\mathbb{R}^{D\cup\partial D}$.
Physically, the interior potential $\mathbf{u}$ satisfies Kirchhoff’s current law, and mathematically, we express it as the discrete Laplacian operator $\Delta_{\gamma}$, defined by
$(\Delta_{\gamma}\mathbf{u})_{p}=\sum_{q\in\mathcal{N}(p)}\gamma_{qp}(\mathbf{u}_{q}-\mathbf{u}_{p}),\quad p\in D.$
One can thus formulate the Dirichlet problem:
$\left\{\begin{aligned} (\Delta_{\gamma}\mathbf{u})_{p}&=0,\quad p\in D,\\ \mathbf{u}_{p}&=\varphi_{p},\quad p\in\partial D,\end{aligned}\right$. (1.1)
where $\varphi\in\mathbb{R}^{\partial D}$ is the Dirichlet boundary condition. The unique solvability of (1.1) is well known [19] for general graphs. Then one can define an injective solution operator $S_{\gamma}$ by
$S_{\gamma}:\mathbb{R}^{\partial D}$ $\to\mathbb{R}^{D\cup\partial D},\quad\varphi\mapsto\mathbf{u},$ (1.2)
where $\mathbf{u}$ is the unique solution of (1.1) with the Dirichlet condition $\varphi$. Assume that each boundary vertex $p\in\partial D$ is connected to precisely one interior vertex, that is, $\mathcal{N}(p)=\{q\}\subseteq D$. Then the boundary current data can be represented as
$\psi_{p}:=\sum_{s\in\mathcal{N}(p)}\gamma_{sp}(\mathbf{u}_{s}-\mathbf{u}_{p})=\gamma_{pq}(\mathbf{u}_{q}-\mathbf{u}_{p}).$
Formally, one can also define
$D_{\gamma}:\mathbb{R}^{D\cup\partial D}$ $\to\mathbb{R}^{\partial D},\quad\mathbf{u}\mapsto\psi.$ (1.3)
The DtN matrix $\Lambda_{\gamma}$ is then the composition of the two operators:
$\Lambda_{\gamma}=D_{\gamma}\circ S_{\gamma}.$
Then the discrete Calderón problem reads: Given the DtN matrix $\Lambda_{\gamma}$, is it possible to uniquely recover the conductivity $\gamma$? In the literature it is also known as the discrete inverse conductivity problem. In their seminal work [19], Curtis and Morrow established the uniqueness of the solution for the inverse problem in two-dimensional square lattices. This study aims to extend the result of Curtis and Morrow [19] to three and higher dimensional lattice structures.
Now we precisely define the graph $G=(E,D,\partial D)$ in a high-dimensional grid. The set of interior vertices $D$ consists of all integer tuples $(x_{i})_{i=1}^{d}\in\mathbb{Z}^{d}$ such that each coordinate adheres to the constraints $1\leq x_{i}\leq n$. This can be expressed as
$D\coloneqq\{(x_{i})_{i=1}^{d}\mid x_{i}\in\mathbb{Z},\ 1\leq x_{i}\leq n\}.$ (1.4)
The boundary vertex set $\partial D$ is comprised of nodes for which exactly one coordinate is equal to either 0 or $n+1$, while the remaining coordinates range over the set $\{1,\ldots,n\}$:
$\partial D$ $\coloneqq\{p\in\mathbb{Z}^{d}:d(p,D)=\min_{q\in D}\|q-p\|_{\ell^{1}}=1\}.$ (1.5)
The edge set $E$ consists of unordered pairs $\{p,q\}$ where $p=(p_{i})_{i=1}^{d}$ and $q=(q_{i})_{i=1}^{d}$ are vertices in $D\cup\partial D$ that meet the following criteria: (i) $\|p-q\|_{\ell^{1}}=1$; and (ii) The pair does not reside entirely within the boundary, i.e., $\{p,q\}\not\subseteq\partial D$. This defines the multi-dimensional hypercubic lattice graphs:
$G=(E,D,\partial D).$ (1.6)
The main contribution of this work is the following uniqueness result for the discrete Calderón problem on multi-dimensional hypercubic graphs.
Theorem 1.1 (Uniqueness).
For the multi-dimensional grids graph $G$ given in (1.6), the DtN matrix $\Lambda_{\gamma}$ uniquely determines the conductivity $\gamma$ on each edge $E$.
Theorem 1.1 generalizes the result of Curtis and Morrow [19] on the uniqueness of the discrete Calderón problem for two-dimensional square lattice to the multi-dimensional case. The proof of Theorem 1.1 employs mathematical induction and a slicing technique that effectively breaks down the complex lattice into more manageable components. The uniqueness of the conductivity $\gamma$ is established through an inductive procedure that is applied slice by slice, based on the slice decomposition strategy. The inductive step based on refined characterization of the algebraic properties of the solution spaces $\mathcal{U}^{(t)}$ on hypercubic lattices. Moreover, the analysis strategy is constructive, and lends itself to an efficient algebraic reconstruction algorithm that can be used to recover the conductivity $\gamma$ from the DtN matrix $\Lambda_{\gamma}$. In Section 5, we present the reconstruction algorithm and numerical illustrations to show the feasibility of the reconstruction.
Last we discuss related works on the discrete Calderón problem. The problem in the two-dimensional case has been extensively studied since the seminal work of Curtis and Morrow [19]. There are many interesting theoretical results for the discrete Calderón problem on planar graphs, including square lattices [19, 20] and circular planar graphs [17, 15, 18, 22]. The work [14] proved a uniqueness result from one boundary measurement under a certain monotonicity assumption. Boyer et al [11] investigated the recovery of a complex-valued conductivity (or complex-valued potential) from the DtN matrix $\Lambda_{\gamma}$, and established the unique recovery up to a set of measure zero.
In sharp contrast, the discrete Calderón problem in the multi-dimensional case has only been scarcely studied so far. Lam and Pylyavskyy [23] investigated the recovery of the conductivity from the so-called $R$-response matrix, proved that the problem generally does not have uniqueness, and established uniqueness for a certain class of purely cylindrical networks. The analysis in [23] relies heavily on grove combinatorics. Araúz et al [1] treated cubic lattices for a particular excitation, and proved the unique recovery of the conductivity from the Robin-to-Neumann map. The proof relies on proving that certain overdetermined boundary value problems on graphs admit a unique solution. Compared with the work [1], we employ the space of admissible excitations as the main tool, and the approach extends readily to general multi-dimensional hypercubic lattice graphs and possibly more general shaped graphs. Moreover, the proposed approach is constructive in the sense that it lends itself directly to an algebraic reconstruction algorithm similar to Curtis-Morrow algorithm for square lattices [20].
The works [24, 13, 21, 25] investigated a closely related inverse problem of recovering the potential in the discrete Schrödinger operator from the (partial) DtN type matrix. Note that due to the lack of a neat Liouville type transformation on graphs, the discrete inverse Schrödinger problem is not equivalent to the discrete Calderón problem. Interestingly, in the work [21], the slicing decomposition was also crucially employed in the uniqueness proof. Although this work shares the spirit with [21], the recovery of the conductivity $\gamma$ requires several additional investigations of the algebraic properties of the solution space $\mathcal{U}^{(t)}$, and thus the analysis is much more intricate. Also we refer interested readers to [4, 3, 16] for other recent important works on inverse problems on graphs.
The rest of the paper is organized as follows. In Section 2, we construct a specifically designed boundary potential that confines nonzero currents to a corner region. In Section 3, we collect three fundamental topological properties of grid graphs, which are essential for the proof of Theorem 1.1. In Section 4, we present the proof of Theorem 1.1. The proof leads to an explicit reconstruction algorithm that incrementally recovers the conductivity $\gamma$, slice by slice, relying solely on the DtN matrix $\Lambda_{\gamma}$. Finally, in Section 5, we present numerical results to illustrate the algorithm derived from the uniqueness proof. Throughout this work, we denote the space of real-valued functions defined on a set $U$ by $\mathbb{R}^{U}$. We extend functions defined on subsets $U\subset V$ to functions on $V$ by assuming they are zero outside $U$ and then we can define the inclusion $\mathbb{R}^{U}\subseteq\mathbb{R}^{V}$. Given a matrix $A:\mathbb{R}^{V_{1}}\rightarrow\mathbb{R}^{V_{2}}$, we define its sub-matrix with column indices $U_{1}\subseteq V_{1}$ and row indices $U_{2}\subseteq V_{2}$ by $A(U_{2};U_{1})$.
2 Corner excitations
In this section, we introduce a special type of boundary excitations that induce well localized potentials and give their important properties. This is the main technical tool to carry out with the inductive steps along the slices. In the proof, the key step is inductive, starting from a subdomain with the known conductivity, to recover the conductivity on a larger region. That is, the domain with known conductivity grows with the induction step, which eventually covers the entire edge set. To rigorously formulate the inductive step, we define a class of growing regions slice by slice as
$L_{t}$ $=\left\{(x_{i})_{i=1}^{d}\in D:\sum_{i=1}^{d}x_{i}=t\right\},$ $L_{t}^{\mathcal{S}}$ $=\bigcup_{\ell=0}^{t}L_{\ell}=\left\{(x_{i})_{i=1}^{d}\in D:\sum_{i=1}^{d}x_{i}\leq t\right\},$ (2.1a)
$K_{t}$ $=\left\{(x_{i})_{i=1}^{d}\in\partial D:\sum_{i=1}^{d}x_{i}=t\right\},$ $K_{t}^{\mathcal{S}}$ $=\bigcup_{\ell=0}^{t}K_{\ell}=\left\{(x_{i})_{i=1}^{d}\in\partial D:\sum_{i=1}^{d}x_{i}\leq t\right\}.$ (2.1b)
The inductive step is that, given the conductivity $\gamma$ on $E(L_{t}^{\mathcal{S}}\cup K_{t}^{\mathcal{S}},L_{t}^{\mathcal{S}}\cup K_{t}^{\mathcal{S}})$, recover $\gamma$ on $E(L_{t+1}^{\mathcal{S}}\cup K_{t+1}^{\mathcal{S}},L_{t+1}^{\mathcal{S}}\cup K_{t+1}^{\mathcal{S}})$. The analysis requires also considering a subset $J_{t}^{\mathcal{S}}$ on the boundary that lies in the lower part of the graph $G$ divided by $L_{t+1}$:
$K_{t}^{+}$ $=\{(x_{i})_{i}\in K_{t}:\max_{i}x_{i}=n+1\},$ $K_{t}^{\mathcal{S}+}$ $=\bigcup_{\ell=0}^{t}K_{\ell}^{+},$ (2.2a)
$K_{t}^{-}$ $=\{(x_{i})_{i}\in K_{t}:\min_{i}x_{i}=0\},$ $K_{t}^{\mathcal{S}-}$ $=\bigcup_{\ell=0}^{t}K_{\ell}^{-}.$ (2.2b)
$J_{t}$ $=K_{t}^{-}\cup K_{t+1}^{+},$ $J_{t}^{\mathcal{S}}$ $=K_{t}^{\mathcal{S}-}\cup K_{t+1}^{\mathcal{S}+}.$ (2.2c)
We refer to Fig. 1 for a schematic illustration of the interface partition in the three-dimensional case.
Refer to caption Refer to caption
$t\leq n+d-1$ t>n+d1t>n+d-1
Figure 1: The interface partition within the domain $D\subset\mathbb{Z}^{3}$. The blue region is in $L_{t+1}^{\mathcal{S}}$.
Intuitively, to recover the conductivity on the current sub-region $L_{t+1}^{\mathcal{S}}\cup K_{t+1}^{\mathcal{S}}$, we select a set of boundary excitations that ”focus” on the sub-region and use its boundary measurements to finish the inductive step. That is, the non-zero current is localized in the sub-region.
$\mathbf{x}$$\mathbf{y}$$\mathbf{z}$$T_{1}^{(t)}$$T_{2}^{(t)}$$T^{(t)}$
Figure 2: $J_{t}^{\mathcal{S}}$ (blue), $L_{t+1}$ (green) and $\partial D\setminus J_{t}^{\mathcal{S}}$ (yellow) in the three dimensional case and its projected view.
The excitations are exactly in the kernel of the following map
$T^{(t)}:=\Lambda_{\gamma}(\partial D\setminus J_{t}^{\mathcal{S}};J_{t}^{\mathcal{S}}).$ (2.3)
We can factorize the map $T^{(t)}$ into two parts: one is the map from the potential supported at the boundary $J_{t}^{\cal S}$ to the potential supported on the boundary $L_{t+1}$, and the other is the map from the potential supported on $L_{t+1}$ to the boundary currents at $\partial D\setminus J_{t}^{\cal S}$. We refer readers to Fig. 2 for a schematic illustration in the three-dimensional case: the sought-for excitations are precisely supported on the blues nodes with the boundary current vanishing at yellows nodes. More precisely, the map $T_{t}^{(1)}$ is defined by
$T_{1}^{(t)}:=S_{\gamma}(L_{t+1};J_{t}^{\mathcal{S}}).$ (2.4)
For $T_{t}^{(2)}$, we first introduce the upper subgraph subgraph $G^{(t)}$ above the interface $L_{t+1}$:
$G^{(t)}=\bigl(E^{(t)},\,D\setminus L_{t+1}^{\mathcal{S}},\,(L_{t+1}\cup\partial D)\setminus J_{t}^{\mathcal{S}}\bigr),$ (2.5)
where the set $E^{(t)}$ contains all edges in the set $E$ connecting nodes within $D\setminus L_{t}^{\mathcal{S}}$. The unique solvability of discrete Dirichlet problem on graph $G^{(t)}$ holds, in view of [19, Remark 3] and hence the DtN map $\Lambda^{(t)}$ on subgraph $G^{(t)}$ is well defined. We can now describe the map $T_{2}^{(t)}$, from the potential at $L_{t+1}$ to the current at $\partial D\setminus J_{t}^{\cal S}$, by a submatrix of $\Lambda^{(t)}$:
$T_{2}^{(t)}:=\Lambda^{(t)}(\partial D\setminus J_{t}^{\mathcal{S}};L_{t+1}):\mathbb{R}^{L_{t+1}}\to\mathbb{R}^{\partial D\setminus J_{t}^{\mathcal{S}}}.$ (2.6)
The next result rigorously establish the factorization.
Proposition 2.1.
For any $t$ in the range $d-1\leq t\leq dn-1$, the operator $T^{(t)}$ can be factorized as $T^{(t)}=T_{2}^{(t)}\circ T_{1}^{(t)}.$ (2.7)
Proof.
Fix any arbitrary $\varphi\in\mathbb{R}^{J_{t}^{\mathcal{S}}}$, and consider the zero extension of $\varphi$ to $\partial D$, still denoted by $\varphi$. The corresponding interior potential is given by $\mathbf{u}:=S_{\gamma}\varphi$. By its definition, $\mathbf{u}$ is $\gamma$-harmonic on $G$, and also within the subgraph $G^{(t)}$. By the definition of the operator $T_{1}^{(t)}$, we have
$\mathbf{u}|_{L_{t+1}}=T_{1}^{(t)}\varphi.$
Since the discrete Dirichlet problem on $G^{(t)}$ has a unique solution, the restriction of $\mathbf{u}$ to $G^{(t)}$ can be represented by
$\mathbf{u}=S_{\gamma}^{(t)}\bigl(T_{1}^{(t)}\varphi\bigr)\quad\text{in }G^{(t)},$
where $S_{\gamma}^{(t)}$ denotes the solution operator on $G^{(t)}$. Now we compute the boundary current and arrive at
$T^{(t)}\varphi=T_{2}^{(t)}\bigl(T_{1}^{(t)}\varphi\bigr).$
Since this relationship holds for any choice of $\varphi\in\mathbb{R}^{J_{t}^{\mathcal{S}}}$, the desired factorization follows. ∎
Remark 2.1.
The factorization in Proposition 2.1 leads directly to the kernel inclusion:
$\ker T_{1}^{(t)}\subseteq\ker T^{(t)}.$ (2.8)
The identity $\ker T_{1}^{(t)}=\ker T^{(t)}$ then follows from injectivity of $T_{2}^{(t)}$, which will be proved in Section 3.
In fact, at the slice $t$, we impose boundary potentials from the kernel of $T_{1}^{(t)}$, record the corresponding boundary currents, and then recover the conductivity between the layers $L_{t}\cup K_{t}$ and $L_{t+1}\cup K_{t+1}$. Since the boundary potentials are chosen from $\ker T_{1}^{(t)}$, the interior potentials are localized within the set $J_{t}^{\mathcal{S}}\cup L_{t}^{\mathcal{S}}$, by the definition of the mapping $T_{1}^{(t)}$. We now define
$\mathcal{U}^{(t)}:=\bigl\{\,S_{\gamma}\varphi\mid\varphi\in\ker T_{1}^{(t)}\bigr\}\big|_{J_{t}^{\mathcal{S}}\cup L_{t}^{\mathcal{S}}},$ (2.9)
where $S_{\gamma}$ is the solution operator defined in (1.2), and the notation $\big|_{J_{t}^{\mathcal{S}}\cup L_{t}^{\mathcal{S}}}$ denotes the restriction of functions from the domain $D$ to the region $J_{t}^{\mathcal{S}}\cup L_{t}^{\mathcal{S}}$, on which the elements $\mathbf{u}\in\mathcal{U}^{(t)}$ are naturally supported. Note that, when recovering the conductivity between $L_{t}\cup K_{t}$ and $L_{t+1}\cup K_{t+1}$, only the boundary potentials in the quotient space
$\ker T_{1}^{(t)}\big/\ker T_{1}^{(t-1)}$
are relevant for the reconstruction subproblem at step $t$. The quotient makes sense in view of the following result of inclusion in the sense $\mathbb{R}^{J_{t-1}^{\mathcal{S}}}\subseteq\mathbb{R}^{J_{t}^{\mathcal{S}}}$.
Proposition 2.2.
For $d\leq t\leq dn-1$:
$\ker T_{1}^{(t-1)}\subseteq\ker T_{1}^{(t)}\$
Proof.
For $\varphi\in\ker T_{1}^{(t-1)}$, let $\mathbf{u}:=S_{\gamma}\varphi$. By construction,
$\mathbf{u}=0\quad\text{in }D\setminus L_{t}\supseteq L_{t+1},$
with $\mathrm{supp}(\varphi)\subseteq J_{t-1}^{\mathcal{S}}\subseteq J_{t}^{\mathcal{S}}$. This implies $\varphi\in\ker T_{1}^{(t)}$ and hence the desired assertion $\ker T_{1}^{(t-1)}\subseteq\ker T_{1}^{(t)}$ holds. ∎
The monotonicity of $\mathcal{U}^{(t)}$ follows directly:
$\mathcal{U}^{(t-1)}=\{S_{\gamma}\varphi:\varphi\in\ker T_{1}^{(t-1)}\}\subseteq\{S_{\gamma}\varphi:\varphi\in\ker T_{1}^{(t)}\}=\mathcal{U}^{(t)}.$
The set of localized potentials $\mathcal{U}^{(t)}$ is a linear subspace of $\mathbb{R}^{J_{t}^{\mathcal{S}}\cup L_{t}^{\mathcal{S}}}$. The following result gives its complement. The following lemma is simply a rewriting of the governing equation in (1.1) for localized solution in $\mathcal{U}^{(t)}$.
Lemma 2.1.
[…]
Proof.
Fix any $\mathbf{u}\in\mathcal{U}^{(t)}$ and extend it onto $D\cup\partial D$ by zero (still denoted by $\mathbf{u}$). By the definition of potential solutions, for $p\in L_{t+1}^{\mathcal{S}}$, we have that $\sum_{q\in\mathcal{N}(p)}\gamma_{pq}(\mathbf{u}_{q}-\mathbf{u}_{p})=0$, or equivalently, that $\mathbf{v}_{p}\cdot\mathbf{u}=0$. Since $\operatorname{supp}(\mathbf{u})=L_{t}^{\mathcal{S}}\cup J_{t}^{\mathcal{S}}$, we may have $\mathbf{v}_{p}\big|_{L_{t}^{\mathcal{S}}\cup J_{t}^{\mathcal{S}}}\cdot\mathbf{u}=0$. Subsequently, we have
$\operatorname{span}\left\{\mathbf{v}_{p}\big|_{L_{t}^{\mathcal{S}}\cup J_{t}^{\mathcal{S}}}:p\in L_{t+1}^{\mathcal{S}}\right\}\subseteq\left(\mathcal{U}^{(t)}\right)^{\perp}.$
Conversely, suppose $\mathbf{u}\in\left(\operatorname{span}_{p\in L_{t+1}^{\mathcal{S}}}\left\{\mathbf{v}_{p}\big|_{L_{t}^{\mathcal{S}}\cup J_{t}^{\mathcal{S}}}\right\}\right)^{\perp}$. Then, for all $p\in L_{t+1}^{\mathcal{S}}$,
$\mathbf{v}_{p}\big|_{L_{t}^{\mathcal{S}}\cup J_{t}^{\mathcal{S}}}\cdot\mathbf{u}=0,$
After extending $\mathbf{u}$ onto $D\cup\partial D$ by zero, the above equations imply that, conversely, we have $\Delta_{\gamma}\mathbf{u}_{p}=0$ for all $p\in D$, where $\Delta_{\gamma}$ is the graph Laplacian. Therefore,
$\mathbf{u}=S_{\gamma}(\mathbf{u}|_{\partial D})\in\mathcal{U}^{(t)},$
and so
$\mathcal{U}^{(t)}\subseteq\left(\operatorname{span}\left\{\mathbf{v}_{p}\big|_{L_{t}^{\mathcal{S}}\cup J_{t}^{\mathcal{S}}}:p\in L_{t+1}^{\mathcal{S}}\right\}\right)^{\perp}.$
Combining both inclusions yields the desired orthogonal decomposition. ∎
Remark 2.2.
The definition of excitations and their properties do not rely on the explicit grid graph structure. Whenever one can divide the graph with boundary nodes into two connected parts, one can always define such localized space, and the respective properties remain valid.
3 Properties of lattice graphs
Note that the unique determination of the conductivity $\gamma$ from the DtN matrix $\Lambda$ is not necessarily valid for a general graph structure. In fact, under the given definitions of $D$ and $\partial D$, we choose the edge set to be complete in $D$ – that is, for every different $p$ and $q$ in $D$, there exists an edge $pq$ in $E$ – then the number of undetermined conductivity parameters, scaling as $O(n^{2d})$, surpasses the size of DtN matrix, which grows as $(2dn^{d-1})^{2}$, for large $n$. This naturally precludes the possibility of unique determination for graph structures, which is characterized by excessive connectivity (i.e. too many edge connections). Thus, the proof of the uniqueness must contain results based on the lattice graph structure.
In this section, we identify several fundamental properties of lattice graphs that may ensure the unique determination of the conductivity $\gamma$. For each integer $t$, consider the subgraph
$G^{\prime(t)}=\bigl(E^{\prime(t)},\,L_{t-1}^{\mathcal{S}},\,J_{t-1}^{\mathcal{S}}\cup L_{t}\bigr),$
where the edge set $E^{\prime(t)}$ consists of all edges connecting nodes in $L_{t-1}^{\mathcal{S}}$ to nodes in $\bar{D}$. Associated to $G^{\prime(t)}$ is the DtN map $\Lambda_{\gamma}^{\prime}$, acting on the boundary potentials. Then we define the operator
$T_{2}^{\prime(t)}:=\Lambda_{\gamma}^{\prime}(J_{t-1}^{\mathcal{S}};L_{t}),$ (3.1)
which maps potentials prescribed on $L_{t}$ to the induced currents on $J_{t-1}^{\mathcal{S}}$.
The next lemma provides the key tool, i.e. to control the potential in the interior domain $L_{t}^{\mathcal{S}}$, using the potential and current taken on $J_{t-1}^{\mathcal{S}}$.
Lemma 3.1.
Let $d\leq t\leq dn$ and suppose $\gamma>0$ is a conductivity on the lattice subgraph $G^{\prime(t)}$. The mixed boundary condition problem for $\mathbf{u}\in\mathbb{R}^{L_{t}^{\mathcal{S}}}$,
$\left\{\begin{aligned} \Delta_{\gamma}\mathbf{u}&=\mathbf{0}\quad\text{in }L_{t-1}^{\mathcal{S}},\\ \mathbf{u}&=\mathbf{0}\quad\text{on }J_{t-1}^{\mathcal{S}},\\ D_{\gamma}\mathbf{u}&=\mathbf{0}\quad\text{on }J_{t-1}^{\mathcal{S}},\end{aligned}\right$. (3.2)
has only trivial solution, i.e., $\mathbf{u}=\mathbf{0}$.
Proof.
We prove the assertion $\mathbf{u}=\mathbf{0}$ in $L_{t}^{\mathcal{S}}$ by mathematical induction. To this end, let
$H_{l}=(J_{t-1}^{\mathcal{S}}\cup L_{t}^{\mathcal{S}})\cap\{(x_{i})_{i}:x_{1}=l\},\quad H_{l}^{\mathcal{S}}=\bigcup_{i=1}^{l}H_{i}.$
Let $\mathbf{e}_{i}=(0,\ldots,0,{1},0,\ldots,0)\in\mathbb{R}^{d}$ be the $j$-th Cartesian coordinate vector. Note that for each $p\in H_{l}\setminus\partial D$ with $l\geq 1$, the node $q=p-\mathbf{e}_{1}$ is connected to exactly one node that lies outside $H_{l-1}^{\mathcal{S}}$, namely $p$ itself. The base case $\mathbf{u}|_{H_{0}}=\mathbf{0}$ holds by assumption. Now assume that $\mathbf{u}|_{H_{l-1}}=\mathbf{0}$. For each $p\in H_{l}\setminus\partial D$, let $q=p-\mathbf{e}_{1}$. Then there holds
$0=\sum_{s\in\mathcal{N}(q)}\gamma_{sq}(u_{s}-u_{q})=\gamma_{pq}u_{p},$
which directly implies $u_{p}=0$. Since $u_{p}=0$ for $p\in H_{l}\cap\partial D$ by assumption, the induction step is complete. Hence, $\mathbf{u}=\mathbf{0}$ in $L_{t}^{\mathcal{S}}$. ∎
The validity of Lemma (3.1) relies on the graph structure. However, the next lemma is independent of the graph structure.
Lemma 3.2.
For every integer $t$ with $d\leq t\leq dn$ and any strictly positive conductivity $\gamma>0$, the following statements are equivalent:
  1. (i)
    Lemma 3.1 holds.
  2. (ii)
    The set of vectors $\{\mathbf{v}_{p}|_{L_{t-1}^{\mathcal{S}}\cup J_{t-1}^{\mathcal{S}}}:p\in L_{t}^{\mathcal{S}}\}$ is linearly independent, where the vectors $\mathbf{v}_{p}$ are defined in (2.10).
  3. (iii)
    The operator $T_{2}^{\prime(t)}$, defined in (3.1), is injective.
Proof.
(i) $\Leftrightarrow$ (ii): The statement (ii) asserts that
$\sum_{p\in L_{t}^{\mathcal{S}}}c_{p}\mathbf{v}_{p}\big|_{L_{t-1}^{\mathcal{S}}\cup J_{t-1}^{\mathcal{S}}}=\mathbf{0}\quad\implies\quad\mathbf{c}=\mathbf{0}.$ (3.3)
By interpreting the vector $\mathbf{c}$ as a potential supported on $G^{\prime(t)}$ with the zero boundary condition $\mathbf{c}|_{J_{t-1}^{\mathcal{S}}}=\mathbf{0}$, the relation (3.3) is equivalent to the uniqueness in (i):
$\left\{\begin{aligned} \Delta_{\gamma}\mathbf{c}&=\mathbf{0}\quad\text{in }L_{t-1}^{\mathcal{S}},\\ \mathbf{c}&=\mathbf{0}\quad\text{on }J_{t-1}^{\mathcal{S}},\\ D_{\gamma}\mathbf{c}&=\mathbf{0}\quad\text{on }J_{t-1}^{\mathcal{S}},\end{aligned}\right.\quad\implies\quad\mathbf{c}=\mathbf{0}.$
(i) $\Leftrightarrow$ (iii): Now suppose that the statement (iii) holds. Consider the problem
$\left\{\begin{aligned} \Delta_{\gamma}\mathbf{u}&=\mathbf{0}\quad\text{in }L_{t-1}^{\mathcal{S}},\\ \mathbf{u}&=\mathbf{0}\quad\text{on }J_{t-1}^{\mathcal{S}},\\ D_{\gamma}\mathbf{u}&=\mathbf{0}\quad\text{on }J_{t-1}^{\mathcal{S}}.\end{aligned}\right$.
Since $D_{\gamma}\mathbf{u}=\mathbf{0}$ on $J_{t-1}^{\mathcal{S}}$, it follows from the injectivity of $T_{2}^{\prime(t)}$ that $\mathbf{u}=\mathbf{0}$ on $L_{t}$. Hence, $\mathbf{u}$ vanishes on the entire boundary of the subgraph $G^{\prime(t)}$. By the uniqueness of the discrete Dirichlet problem on $G^{\prime(t)}$, we conclude $\mathbf{u}=\mathbf{0}$ on $L_{t}^{\mathcal{S}}$. Conversely, suppose that the statement (i) holds. If $T_{2}^{\prime(t)}\mathbf{x}=\mathbf{0}$, then the solution $\mathbf{u}$ to
$\left\{\begin{aligned} \Delta_{\gamma}\mathbf{u}&=\mathbf{0}\quad\text{in }L_{t-1}^{\mathcal{S}},\\ \mathbf{u}&=\mathbf{0}\quad\text{on }J_{t-1}^{\mathcal{S}},\\ \mathbf{u}&=\mathbf{x}\quad\text{on }L_{t},\end{aligned}\right$.
also satisfies
$\left\{\begin{aligned} \Delta_{\gamma}\mathbf{u}&=\mathbf{0}\quad\text{in }L_{t-1}^{\mathcal{S}},\\ \mathbf{u}&=\mathbf{0}\quad\text{on }J_{t-1}^{\mathcal{S}},\\ D_{\gamma}\mathbf{u}&={T_{2}^{\prime(t)}\mathbf{x}}=\mathbf{0}\quad\text{on }J_{t-1}^{\mathcal{S}}.\end{aligned}\right$.
By (i), this implies $\mathbf{u}=\mathbf{0}$ on $L_{t}^{\mathcal{S}}$, and thus $\mathbf{x}=\mathbf{0}$. ∎
Now we focus on lattice graphs and provide several key properties. The first one summarizes the preceding arguments.
Proposition 3.1 (Property 1: Injectivity).
For every $t$, the operators $T_{2}^{(t)}$ and $T_{2}^{\prime(t)}$ are injective.
Proof.
The injectivity of $T_{2}^{\prime(t)}$ follows directly from Lemmas 3.2 and 3.1. By symmetry, the operator $T_{2}^{(t)}$ is also injective. ∎
Next we introduce a slightly modified graph relative to the corner subgraph $G^{\prime(t)}$, from which we remove one node $p$ on the boundary $L_{t}$. Specifically, we define
$G^{\prime\prime(t)}_{p}:=\bigl(E_{p}^{\prime\prime(t)},\;L_{t-1}^{\mathcal{S}}\setminus M_{p},\;(J_{t-1}^{\mathcal{S}}\cup L_{t}\cup M_{p})\setminus\{p\}\bigr),$ (3.4)
where the set $M_{p}:=\mathcal{N}(p)\cap\bigl(J_{t-1}^{\mathcal{S}}\cup L_{t-1}^{\mathcal{S}}\bigr)$ is the neighbours of $p$ within the preceding layer, and the new edge set $E_{p}^{\prime\prime(t)}$ is given by
$E_{p}^{\prime\prime(t)}:=E^{\prime(t)}\setminus\{pq:q\in M_{p}\}.$
Note that if $p$ is adjacent to the boundary, isolated boundary nodes may arise. The boundary current on such nodes is set to zero. Fig. 3 gives a schematic illustration of the removal of the node $p$ on the slice $t$.
$\times$$\times$$\times$$\times$
Figure 3: Node deletion for dimension $d=3$. $L_{t}$ (blue), $L_{t+1}$ (black) and $K_{t}^{-}$ (orange) for $t\leq n+d-1$.
The following lemma establishes the unique recovery of the potential on the reduced vertex set $L_{t}^{\mathcal{S}}\setminus\{p\}$ from the boundary measurements on $J_{t-1}^{\mathcal{S}}$, i.e., the uniqueness of interior potential recovery on the grids $G^{\prime\prime(t)}$.
Lemma 3.3 (Property 2: Uniqueness on the reduced subgraph).
Let $d\leq t\leq dn$ and fix any node $p\in L_{t}$. Suppose the conductivity $\gamma>0$ is distributed on the subgraph $G_{p}^{\prime\prime(t)}$ defined in (3.4). Consider the boundary value problem:
$\left\{\begin{aligned} \Delta_{\gamma}\mathbf{u}&=\mathbf{0}\quad\text{in }L_{t-1}^{\mathcal{S}}\setminus M_{p},\\ \mathbf{u}&=\mathbf{0}\quad\text{on }J_{t-1}^{\mathcal{S}},\\ D_{\gamma}\mathbf{u}&=\mathbf{0}\quad\text{on }J_{t-1}^{\mathcal{S}},\end{aligned}\right$. (3.5)
the only solution $\mathbf{u}$ supported on $L_{t}^{\mathcal{S}}\setminus\{p\}$ is the trivial solution $\mathbf{u}=\mathbf{0}$.
Proof.
Let the deleted node $p$ be located at $(p_{i})_{i}$. We can recover the potential $\mathbf{u}$ in the region $\{(x_{i})\in D:x_{1}<p_{1}\}$ using the same argument in Lemma 3.1. By changing the inductive direction from $\mathbf{e}_{1}$ to $\mathbf{e}_{j}$ with $j>1$, $\mathbf{u}$ can also be recovered in the region
$\{(x_{i})_{i}:\exists i\in\{1,\ldots,d\}\text{ such that }x_{i}<p_{i}\},$
which together with $L_{t}^{\mathcal{S}}\setminus\{p\}$ completes the recovery. ∎
This result yields an important relation among the vectors $\mathbf{v}_{q}$ associated with nodes in the layers $L_{t}^{\mathcal{S}}$, given in the following corollary.
Corollary 3.2.
The vectors $\mathbf{v}_{q}$ satisfy the following properties:
  1. (i)
    The collection of vectors $\{\mathbf{v}_{q}|_{L_{t-1}^{\mathcal{S}}\setminus M_{p}}\}_{q\in L_{t}^{\mathcal{S}}\setminus\{p\}}$ is linearly independent.
  2. (ii)
    If $\mathbf{w}\in\operatorname{span}\{\mathbf{v}_{q}|_{L_{t-1}^{\mathcal{S}}}\}_{q\in L_{t}^{\mathcal{S}}}$ is supported on $M_{p}$, then there exists $\alpha\in\mathbb{R}$ such that
    $\mathbf{w}=\alpha\mathbf{v}_{p}|_{L_{t-1}^{\mathcal{S}}}.$
Proof.
The assertion (i) follows from the argument for the equivalence (i) $\Leftrightarrow$ (ii) in Lemmas 3.2 and $3.$ To see assertion (ii), suppose that $\mathbf{w}=\sum_{q\in L_{t}^{\mathcal{S}}}c_{q}\mathbf{v}_{q}|_{L_{t-1}^{\mathcal{S}}}$ is supported on $M_{p}$. By restricting $\mathbf{w}$ to $L_{t-1}^{\mathcal{S}}\setminus M_{p}$, we have
$\mathbf{0}=\sum_{q\in L_{t}^{\mathcal{S}}\setminus\{p\}}c_{q}\mathbf{v}_{q}|_{L_{t-1}^{\mathcal{S}}\setminus M_{p}}.$
By part (i), it follows that all coefficients $c_{q}=0$ for $q\neq p$. Hence, $\mathbf{w}=\alpha\mathbf{v}_{p}|_{L_{t-1}^{\mathcal{S}}}$ for some $\alpha\in\mathbb{R}$. ∎
Finally, we state an elementary observation on connectivity:
Lemma 3.4 (Property 3: Interface connectivity).
The neighborhood $\mathcal{N}(L_{t}\cup K_{t})$ of the set $L_{t}\cup K_{t}$ satisfies
$\mathcal{N}(L_{t}\cup K_{t})\subseteq L_{t-1}\cup K_{t-1}\cup L_{t+1}\cup K_{t+1}.$
Notably, there are no edges connecting nodes within $L_{t}\cup K_{t}$ itself.
4 Proof of Theorem 1.1
We now prove Theorem 1.1, i.e., the unique determination of $\gamma$ from the DtN map $\Lambda_{\gamma}$. The proof is based on mathematical induction slice by slice.
4.1 Extract the solution space from DtN data
The only available data for the recovery oof the conductivity, aside from the graph structure, is the DtN matrix $\Lambda_{\gamma}$. In order to recover the conductivity using the boundary measurements from solutions in $\mathcal{U}^{(t)}$, we shall first prove that the respective boundary potentials $\ker T_{1}^{(t)}$ can be directly determined by $\Lambda_{\gamma}$. In fact, $T^{(t)}=\Lambda_{\gamma}(\partial D\setminus J_{t}^{\mathcal{S}};J_{t}^{\mathcal{S}})$ is a submatrix of $\Lambda_{\gamma}$, which achieves the unique determination by combining the following lemma:
Lemma 4.1.
For all $d-1\leq t\leq dn-1$, we have
$\ker T_{1}^{(t)}=\ker T^{(t)}.$
Proof.
By the factorization identity
$T^{(t)}=T_{2}^{(t)}\circ T_{1}^{(t)},$
in Proposition 2.1, for any $\varphi\in\ker T_{1}^{(t)}$ we have
$T^{(t)}\varphi=T_{2}^{(t)}(T_{1}^{(t)}\varphi)=T_{2}^{(t)}(\mathbf{0})=\mathbf{0}.$
This shows the inclusion
$\ker T_{1}^{(t)}\subseteq\ker T^{(t)}.$
Conversely, Proposition 3.1 guarantees that $T_{2}^{(t)}$ is injective. Therefore, for any $\varphi\in\ker T^{(t)}$, we have
$\mathbf{0}=T^{(t)}\varphi=T_{2}^{(t)}(T_{1}^{(t)}\varphi),$
which implies $T_{1}^{(t)}\varphi=\mathbf{0}$. Hence, $\ker T^{(t)}\subseteq\ker T_{1}^{(t)}$. Combining these two inclusions yields the desired equality $\ker T_{1}^{(t)}=\ker T^{(t)}$. ∎
4.2 Slice-by-slice reconstruction of conductivity
The base case begins with the layer that is closest to the origin. Note that $L_{d}=\{(1,1,\cdots,1)\}$ and $K_{d-1}=\{(1,\cdots,1,\underbrace{0}_{\text{the i th index}},1,\cdots,1):i=1,\cdots,d\}$. We begin the reconstruction of the conductivity $\gamma$ by focusing on the subset of edges connecting the node sets $K_{d-1}$ and $L_{d}$:
$E(K_{d-1},L_{d}):=\{\,pq\in E:p\in K_{d-1},\ q\in L_{d}\,\}.$
Proof – Recovery of $\gamma$ on $E(K_{d-1},L_{d})$.
By Lemma 2.1, the orthogonal complement of
$\mathcal{U}^{(d-1)}|_{K_{d-1}}=\ker T_{1}^{(d-1)}$
is one-dimensional. This complement is spanned by a vector that encodes precisely the conductivity values
$(\gamma_{pq})_{q\in K_{d-1}},\quad\text{with }p=(1,\ldots,1)\in L_{d}.$
Thus the subspace $\ker T_{1}^{(d-1)}$ uniquely determines the conductivity values $\gamma$ on the edge set $E(K_{d-1},L_{d})$. ∎
Next we decompose the edge set $E^{t}$ as
$E^{t}:=\bigcup_{i=d-1}^{t}E_{i},$ (4.1)
where for each $i\geq d-1$, the subset
$E_{i}:=E(K_{i}^{-},L_{i+1})\cup E(L_{i},K_{i+1}^{+})\cup E(L_{i},L_{i+1})$ (4.2)
collects all edges connecting nodes in $K_{i}\cup L_{i}$ to those in $K_{i+1}\cup L_{i+1}$. The interface graph
$G_{t}:=\bigl(L_{t}\cup J_{t}\cup L_{t+1},\ E_{t}\bigr)$ (4.3)
thus encodes the connectivity structure relevant to the layers; see Fig. 4 for a schematic illustration on cubic lattices.
Figure 4: $L_{t}$ (blue), $L_{t+1}$ (black), $K_{t}^{-}$ (orange) and $K_{t+1}^{+}$ (green) for $t\leq n+d-1$ (left) and t>n+d1t>n+d-1 (right) in the 3D case.
Inductive proof of Theorem 1.1: Suppose that the conductivity $\gamma$ on the edges $E^{t-1}$ is known (this holds true for the base case $t=d$). We impose the boundary potential by $\varphi\in\ker T_{1}^{(t)}$, and record the resulting boundary current $\Lambda_{\gamma}\varphi$. Given the conductivity on $E^{t-1}$, we can determine the interior potential distribution $\mathbf{u}$, which is supported within $L_{t}^{\mathcal{S}}$, where the existence of a unique solution is ensured by Lemma 3.1 . Then $\gamma|_{E_{t}}$ should satisfy ∑q∈N(p)γpq(up-uq)=0, for p∈Lt∪Lt+1,∑q∈N(p)γpq(up-uq)=ψp,  for p∈Jt. These relations can be expressed as a system of linear equations of the only unknown $\bm{\gamma}$:
$\mathbf{J}_{p}^{\mathbf{u}}\cdot\bm{\gamma}=C_{p},\text{ for all }p\in L_{t}\cup L_{t+1}\cup J_{t}{\text{ and }\mathbf{u}\in\mathcal{U}^{(t)},}$ (4.4)
where for each edge $e=pq$ in $E_{t}$, the vector $\mathbf{J}_{p}^{\mathbf{u}}\in\mathbb{R}^{E_{t}}$ is defined component-wise by
$\mathbf{J}_{p}^{\mathbf{u}}(e)=\begin{cases}\mathbf{u}_{p}-\mathbf{u}_{q},&\text{if }e=pq,\\ 0,&\text{otherwise}.\end{cases}$ (4.5)
The existence of a solution is always valid, and the solutions necessarily contain the true conductivity that generate the data, i.e., the DtN matrix $\Lambda_{\gamma}$. Furthermore, the next theorem gives the uniqueness of the solutions and thereby achieves the reconstruction of a single-layer conductivity.
Theorem 4.1.
In $\mathbb{R}^{E_{t}}$, there holds
$\operatorname{span}\left(\left\{\mathbf{J}_{p}^{\mathbf{u}}:\mathbf{u}\in\mathcal{U}^{(t)},p\in L_{t}\cup L_{t+1}\cup J_{t}\right\}\right)^{\perp}=\{0\}.$
Proof.
We first investigate the space $\operatorname{span}\{\mathbf{J}_{p}^{\mathbf{u}}:\mathbf{u}\in\mathcal{U}^{(t)}\}$ for each node $p$ under three cases:
Figure 5: Illustration of $\mathbf{J}_{p}^{\mathbf{u}}$ with different $p$ in the 3D case with $L_{t}$ (blue), $L_{t+1}$ (black), $K_{t}^{-}$ (orange) and $K_{t+1}^{+}$ (green) .
Case 1: $p\in L_{t+1}$. Let $M_{p}=(L_{t}\cup K_{t})\cap\mathcal{N}(p)$, where $\mathcal{N}(p)$ is the set of neighbors of $p$. Consider the subspace
$\operatorname{span}\left\{\mathbf{u}|_{M_{p}}:\mathbf{u}\in\mathcal{U}^{(t)}\right\}.$
If $\mathbf{w}\in\mathbb{R}^{M_{p}}$ is orthogonal to the subspace $\operatorname{span}\{\mathbf{u}|_{M_{p}}:\mathbf{u}\in\mathcal{U}^{(t)}\}$, then by Lemma 2.1, $(\mathbf{w},\mathbf{0})$ belongs to $\operatorname{span}\{\mathbf{v}_{q}|_{L_{t}}:q\in L_{t+1}^{\mathcal{S}}\}$ and supports in $M_{p}$. Corollary 3.2 implies $\mathbf{w}=\alpha\mathbf{v}_{p}|_{L_{t}}$ for some $\alpha\in\mathbb{R}$. By Lemma 3.4, the set $M_{p}$ contains exactly the nodes that are connected to $p$ in the graph $G_{t}$. Hence, by noting that $u_{p}=0$, we have
$\left(\operatorname{span}\{\mathbf{J}_{p}^{\mathbf{u}}:\mathbf{u}\in\mathcal{U}^{(t)}\}\right)^{\perp}=\operatorname{span}\{\bm{\gamma}|_{\mathcal{E}(p)}\}\times\mathbb{R}^{E\setminus\mathcal{E}(p)},$ (4.6)
where $\mathcal{E}(p)$ denotes edges incident to $p$, and $\bm{\gamma}|_{\mathcal{E}(p)}$ is supported on the edges directed toward $p$.
Case 2: $p\in L_{t}$ with no connection to $K_{t+1}^{+}$, or $p\in K_{t}^{-}$. Again, by Lemma 3.4, $\mathcal{N}(p)\cap G_{t}=\mathcal{N}(p)\cap L_{t+1}$. Furthermore, we have $\mathbf{u}|_{\mathcal{N}(p)\cap L_{t+1}}=0$. Let $\mathbf{1}_{\mathcal{E}(p)}$ denote the indicator vector on edges incident to $p$. Therefore, to prove
$\operatorname{span}\{\mathbf{J}_{p}^{\mathbf{u}}:\mathbf{u}\in\mathcal{U}^{(t)}\}=\operatorname{span}\{\mathbf{1}_{\mathcal{E}(p)}\},$ (4.7)
it suffices to show that there exists $\mathbf{u}\in\mathcal{U}^{(t)}$ with $\mathbf{u}_{p}\neq 0$. Indeed, if no such $\mathbf{u}$ exists, then the canonical basis vector $\mathbf{e}_{p}\in\mathbb{R}^{J_{t}^{\mathcal{S}}\cup L_{t}^{\mathcal{S}}}$ is orthogonal to $\mathcal{U}^{(t)}$ and supported on some $M_{q}$ with $q\in\mathcal{N}(p)\cap L_{t+1}$. By Corollary 3.2, this implies $\mathbf{e}_{p}=\alpha\mathbf{v}_{q}|_{L_{t}}$ for some $\alpha\in\mathbb{R}$, which contradicts the positivity of $\gamma$.
Case 3: $p\in L_{t}$ connected to $q\in K_{t+1}^{+}$. By an argument analogous to Case 2, one readily verifies that
$\operatorname{span}\{\mathbf{J}_{p}^{\mathbf{u}}:\mathbf{u}\in\mathcal{U}^{(t)}\}+\operatorname{span}\{\mathbf{J}_{q}^{\mathbf{u}}:\mathbf{u}\in\mathcal{U}^{(t)}\}=\underbrace{\mathbb{R}}_{\text{edge }pq}\times\operatorname{span}\{\mathbf{1}_{\mathcal{E}(p)\setminus pq}\}\times\underbrace{\mathbf{0}}_{\text{edges }E\backslash\mathcal{E}(p)},$ (4.8)
where $\mathbf{1}_{\mathcal{E}(p)\setminus pq}$ denotes the indicator vector on edges incident to $p$ excluding the edge $pq$.
Now suppose that $\mathbf{w}$ lies in the orthogonal complement, i.e.,
$\mathbf{w}\in\operatorname{span}\left(\left\{\mathbf{J}_{p}^{\mathbf{u}}:\mathbf{u}\in\mathcal{U}^{(t)},p\in L_{t}\cup L_{t+1}\cup J_{t}\right\}\right)^{\perp}.$ (4.9)
We claim $\mathbf{w}=\mathbf{0}$. First, fix any node $p\in L_{t+1}$. Since $\mathbf{w}$ is orthogonal to each $\mathbf{J}_{p}^{\mathbf{u}}$, its restriction to edges incident to $p$ is proportional to $\bm{\gamma}|_{\mathcal{E}(p)}$ as in (4.6). Hence, there exists $c_{p}\in\mathbb{R}$ such that
$\mathbf{w}|_{\mathcal{E}(p)}=c_{p}\bm{\gamma}|_{\mathcal{E}(p)}.$ (4.10)
Next, the orthogonality condition $\mathbf{w}\perp\mathbf{J}_{p}^{\mathbf{u}}$ for all $p\in L_{t}\cup J_{t}$, together with (4.8) and (4.7), implies the linear constraints
$\sum_{q\in\mathcal{N}(p)\cap L_{t+1}}c_{q}\gamma_{pq}=0,\quad\forall p\in L_{t}\cup K_{t}^{-}.$ (4.11)
We extend $\mathbf{c}$ by setting $\mathbf{c}=0$ on $L_{t}^{\mathcal{S}}\cup J_{t}^{\mathcal{S}}$, so that the boundary nodes $J_{t}^{\mathcal{S}}$ have zero potential and zero current. Then $\mathbf{c}$ solves the boundary value problem
$\left\{\begin{aligned} \Delta_{\gamma}\mathbf{c}&=\mathbf{0}\quad\text{in }L_{t},\\ \mathbf{c}&=0\quad\text{on }J_{t}^{\mathcal{S}},\\ D_{\gamma}\mathbf{c}&=0\quad\text{on }J_{t}^{\mathcal{S}}.\end{aligned}\right$.
By Lemma 3.1, the only solution is $\mathbf{c}=0$ on $L_{t+1}$. This and the identity (4.10) imply that $\mathbf{w}$ vanishes on all edges incident to nodes in $L_{t+1}$, except possibly those incident to nodes in $K_{t+1}^{+}$. Finally, by (4.8), we conclude that $\mathbf{w}=\mathbf{0}$ everywhere. This completes the proof of the theorem. ∎
Theorem 4.2.
The data pairs
$\{(\varphi,\Lambda_{\gamma}\varphi):\varphi\in\ker T_{1}^{(t)}\}$
uniquely determine the conductivity $\bm{\gamma}$ on the entire edge set $E$.
Proof.
The true conductivity $\bm{\gamma}$ lies in the set {γJ_p^uγ=C_p for all p∈L_t∪L_t+1∪J_t,uU^(t)}. If $\bm{\gamma}_{1}$ and $\bm{\gamma}_{2}$ lie in the set, by Theorem 4.1, we have
$\gamma_{1}-\gamma_{2}\in\operatorname{span}\left(\left\{\mathbf{J}_{p}^{\mathbf{u}}:\mathbf{u}\in\mathcal{U}^{(t)},p\in L_{t}\cup L_{t+1}\cup J_{t}\right\}\right)^{\perp}=\{0\},$
which establishes uniqueness. ∎
Upon combining Lemma 4.1 and Theorem 4.2, the proof of Theorem 1.1 is complete. $\blacksquare$
5 Numerical reconstructions and discussions
The uniqueness proof directly lends itself to a slice-by-slice reconstruction algorithm, which can be viewed as a multi-dimensional analogue of the algebraic reconstruction algorithm due to Curtis and Morrow [20] for square lattices. The detailed procedure is listed in Algorithm 1. When recovering the conductivity on $E_{t}$, it employs a basis of the quotient space $\ker T^{(t)}/\ker T^{(t-1)}$. Like Curtis-Morrow algorithm [20], each iteration of the algorithm involves only simple algebraic operations. One essential step per loop is to solve for the potential with the Cauchy data specified on the boundary $J_{t-1}^{\mathrm{S}}$. Similar to the Cauchy problem for the Laplace equation, the problem is expected to be ill-conditioned.
Algorithm 1 The reconstruction algorithm for the conductivity $\bm{\gamma}$
1:DtN map $\Lambda_{\gamma}$, grid size $n$, and dimension $d$.
2:Approximation of the conductivity $\gamma$.
3:Initialize $\gamma\leftarrow\mathbf{0}$.
4:for t=d1t=d-1 to $d\lceil\frac{n+1}{2}\rceil$ do
5:  Compute the kernel $K_{t}\coloneqq\ker\Lambda_{\gamma}(\partial D\setminus J_{t}^{\mathcal{S}},J_{t}^{\mathcal{S}})$.
6:  for each $\varphi_{i}$ spanning $K_{t}/K_{t-1}$ do
7:   Compute $\mathbf{u}_{i}\leftarrow$ Potential($\gamma,\varphi_{i},\Lambda_{\gamma}\varphi_{i},t$).
8:   Evaluate $\mathbf{J}_{p}^{\mathbf{u}_{i}}$ for all $p\in L_{t}\cup L_{t+1}\cup J_{t}$ using (4.5).
9:   Compute $C_{p}$ according to (4.4).
10:  end for
11:  Determine $\gamma|_{E_{t}}$ by solving the system
$\left\{\bm{\gamma}:\mathbf{J}_{p}^{\mathbf{u}}\cdot\bm{\gamma}=C_{p},\quad\forall p\in L_{t}\cup L_{t+1}\cup J_{t},\quad\forall\mathbf{u}=\mathbf{u}_{i}\right\}.$
12:end for
13:function Potential($\gamma,\varphi,\psi,t$)
14:  Find $\mathbf{u}$ satisfying
$\left\{\begin{aligned} \Delta_{\gamma}\mathbf{u}&=\mathbf{0}\quad\text{in }L_{t-1}^{\mathcal{S}},\\ \mathbf{u}&=\varphi\quad\text{on }J_{t-1}^{\mathcal{S}},\\ D_{\gamma}\mathbf{u}&=\psi\quad\text{on }J_{t-1}^{\mathcal{S}}.\end{aligned}\right$.
15:end function
We present numerical results for cubic lattices to illustrate the feasibility of the algorithm. The relevant numerical results are shown in Figs. 6 and 7 for cubic lattices of different sizes $n$. Note that Algorithm 1 initiates the reconstruction process from a single corner of the cubic lattice. However, numerically the error grows as the iteration proceeds. The displayed reconstructions combine the eight solutions obtained by starting from each of the eight corners, and for each edge in the graph, we select the solution corresponding to the closest corner.
The results in Figs. 6 and 7 show that the algorithm indeed can accurately recover the conductivity from the DtN matrix $\Lambda_{\gamma}$ (when the lattice size $n$ is not large), which corroborates the proof strategy. For each fixed $n$, the pointwise error increases towards the center (i.e., the central part has the largest error), exhibiting clearly a depth-dependent resolution: the conductivity on the edges near the corners can be more accurately recovered to recover than that are far away from the corners. Moreover, as the lattice size $n$ increases from 8 to 13, the maximum error increases dramatically, from $10^{-9}$ (for $n=8$) to $10^{0}$ (for $n=13$), and the dominant error occurs in the central part of the cubic lattice. Nonetheless, the reconstruction accuracy near the corners is still quite satisfactory. Note that the error arises purely from the round-off error of the floating point operations in the algorithm. This issue is a clear manifestation of the severe ill-posed nature of the discrete Calderón problem: the problem is exponentially ill-posed with respect to the lattice size $n$, and the phenomenon is well known also in the 2D case [19, 2]. Therefore, in the presence of data noise, to obtain reasonable conductivity reconstructions, incorporating suitable regularization is indispensable.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
$n=8$ $n=9$ $n=10$
Figure 6: The exact conductivity (top) and the logarithmic error $\log_{10}|\mathbf{e}_{\gamma}|$ (bottom) of the recovered conductivity for the three-dimensional cubic lattice.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
$n=11$ $n=12$ $n=13$
Figure 7: The exact conductivity (top) and the logarithmic error $\log_{10}|\mathbf{e}_{\gamma}|$ (bottom) of the recovered conductivity for the three-dimensional cubic lattice.